how would you prove that triangle bca ia a right angled triangle?? co-ordinates on plane are: a(-4,3) b(2,-4) c(0,4)
Answer by oscargut(2103)

You can find distances
d(C,A)=sqrt(4^2+1^2)=sqrt(17)
d(C,B)=sqrt(2^2+8^2)=sqrt(68)
d(A,B)=sqrt(6^2+7^2)=sqrt(85)
then d(A,B)^2=d(C,A)^2+d(C,B)^2 ehen using Pythagoras Theorem
triangle ABC is a right angled triangle in C
(You can also find lines AC and BC and verify they form a right angle)